OpenStudy (anonymous):

solve for y in y+3 = sqrt(3y+7)

OpenStudy (anonymous):

square both sides

OpenStudy (anonymous):

but what do u do when you get the y^2+9 = 3y +7

OpenStudy (anonymous):

that's not what you get when you square (y+3). (y+3)^2 = (y+3)(y+3) = y^2 + 6y + 9. now equate this with the right side of 3y + 7 and solve. Remember to check your solution(s)...

OpenStudy (anonymous):

the answer would be y=-2, and y=-1